Let the independent random variables X1 and X2 be N(0,1) and "\chi" 2(r) respectively. Let Y1=X1/(X2)^(1/2)/r and Y2=X2

i) Find the joint pdf of Y1 and Y2

ii) Determine the marginal pdf of Y1 and show that Y1 has a t distribution


1
Expert's answer
2022-02-01T08:42:14-0500

"X_1\sim N(0,1)" and its probability density function is given as,

"g(x_1)={1\over\sqrt{2\pi}}e^{-{x_1^2\over2}},\space 0\lt x_1\lt\infin"

and,

"X_2\sim \chi^2(r)" and its probability density function is given as,

"h(x_2)={1\over\varGamma({r\over2})2^{r\over2}}x_2^{{r\over2}-1}e^{-{x_2\over2}},\space 0\lt x_2\lt\infin"

Since the random variables "X_1" and "X_2" are independent, their joint probability density function is given as,

"f(x_1,x_2)=g(x_1)\times h(x_2)". So,

"f(x_1,x_2)={1\over\sqrt{2\pi}}e^{-{x_1^2\over2}}{1\over\varGamma({r\over2})2^{r\over2}}x_2^{{r\over2}-1}e^{-{x_2\over2}},\space -\infin\lt x_1\lt \infin,\space 0\lt x_2\lt\infin\\0,\space elsewhere"

Define,

"Y_1={X_1\over\sqrt{(X_2)\over r}}" and "Y_2=X_2". They are one to one mapping of the set, "S=\{(x_1,x_2):-\infin\lt x_1\lt\infin, 0\lt x_2\lt\infin\}" onto the set,

"R=\{(y_1,y_2),-\infin\lt y_1\lt\infin, 0\lt y_2\lt\infin\}"

So,

"X_2=Y_2" and "X_1=Y_1\sqrt{Y_2\over r}={Y_1(Y_2)^{1\over2}\over \sqrt{r}}"

The Jacobian of transformation is,

"J=\begin{vmatrix}
 {d_{x_1}\over d_{y_1}} & {d_{x_1}\over d_{y_2}} \\
 {d_{x_2}\over d_{y_1}} & {d_{x_2}\over d_{y_2}}
\end{vmatrix}=\begin{vmatrix}
 {\sqrt{y_2\over r}} & {y_1\over2\sqrt{r\times y_2}} \\
 0 & 1
\end{vmatrix}=\sqrt{{y_2\over r}}"

Then, the joint pdf of "Y_1" and "Y_2" is given by,

"g(y_1,y_2)=g(y_1\sqrt{y_2\over r},y_2)|J|\\=
{1\over\sqrt{2\pi}\varGamma({r\over2})2^{r\over2}}y_2^{{r\over2}-1}e^{-{{1\over2}y_1^2}{y_2\over r}}e^{-{y_2\over2}}\sqrt{{y_2\over r}}\\=
{1\over\sqrt{2\pi}\varGamma({r\over2})2^{r\over2}}y_2^{{r\over2}-1}e^{-{y_2\over2}(1+{y_1^2\over r})}{y_2^{1\over2}\over\sqrt{r}},\space -\infin\lt y_1\lt\infin,\space 0\lt y_2\lt\infin"

The marginal pdf of "Y_1" is given by,

"g(y_1)=\displaystyle\int^{\infin}_0 g(y_1,y_2)dy_2=\displaystyle\int^{\infin}_0{1\over\sqrt{2\pi}\varGamma({r\over2})2^{r\over2}}y_2^{{{1\over2}(r+1)}-1}e^{-{y_2\over2}(1+{y_1^2\over r})}{y_2^{1\over2}\over\sqrt{r}}dy_2"

We make the substitution,

"w={y_2(1+{y_1^2\over r})\over2}" and "{dw\over dy_2}={1+{y_1^2\over r}\over 2}\implies dw={1+{y_1^2\over r}\over 2}dy_2"

We can now write the integral as,

"\displaystyle\int_0^{\infin}{1\over\sqrt{2\pi r}\varGamma({r\over2})2^{r\over2}}({2w\over1+{y_1^2\over r}})^{{1\over2}(r+1)-1}e^{-w}({2\over1+{y_1^2\over r}})dw"

Next, we introduce "\varGamma({r+1\over2})"

"={1\over\sqrt{2\pi r}\varGamma({r\over2})2^{r\over2}}({2\over1+{y_1^2\over r}})^{{1\over2}(r+1)}\varGamma({r+1\over2})\displaystyle\int_0^{\infin}{w^{{1\over2}(r+1)-1}e^{-w}\over\varGamma({r+1\over2})}dw"

The integral part is equal to 1.0 because it is a pdf of the Gamma distribution with parameters "\alpha={r+1\over2}" and "\beta=1"

So we have,

"{1\over\sqrt{2\pi}\varGamma({r\over2})2^{r\over2}}({2\over1+{y_1^2\over r}})^{{1\over2}(r+1)}\varGamma({r+1\over2})" and we can write as,

"g(y_1)=({\varGamma({r+1\over2})\over\sqrt{\pi r}\varGamma({r\over2})})({1\over(1+{y_1^2\over r})^{r+1\over 2}}),\space -\infin\lt y_1\lt \infin\\0,\space elsewhere"

Thus, if "X_1\sim N(0,1)" and "X_2\sim \chi^2(r)" and "X_1" and "X_2" are independent then, the random variable "Y_1={X_1\over\sqrt{X_2\over r}}" has a t distribution with "r" degrees of freedom.

Therefore,

"Y_1\sim t-distribution(r)."